# Edge AI: A Taxonomy, Systematic Review and Future Directions

### v1

@@ V-A3 Edge
The concept of edge computing is relative to cloud computing. The processing method of cloud computing is to upload all data to the cloud data center or server in the computing resource set for processing. Any request to access this information must be submitted to the cloud for processing [63]. Edge computing is a computing model that moves resources and provides edge intelligent services at the network edge near objects or data sources, by which improve service quality and information security. In brief, edge computing analyzes the data produced by the terminal directly in the local device or network near the data generation, without transmitting the data to the cloud data processing center [47]. Compared with cloud computing, edge computing has shorter network latency, less resource consumption, and higher security of the data to the cloud data processing center.
The so-called edge AI combines two emerging technologies: edge computing and AI. However, the implementation of edge computing is based on the same basic premise, that is, generating, collecting, storing, processing and managing data locally rather than in remote data centers. Edge AI improves this concept to the device level, using ML to imitate human reasoning to reach user interaction points, such as computers, edge servers or IoT devices [48]. Typically, these devices can operate without an Internet connection and make decisions independently. Well-known examples of edge AI technologies include virtual assistants, such as GPT-4o, Apple’s Siri, or Amazon Alexa. When the user says ”hey”, these tools will recognize and learn what the user is saying (i.e. ML), interact with cloud-based application programming interfaces (APIs), and store the learned knowledge locally.
@@ V-C2 Mobile
The emergence of emerging AI such as DL has brought new innovations to mobile animal networking. In the field of mobile IoT, edge AI achieves real-time and efficient data processing by deploying the computing power of AI at the edge of devices. For example, in smart health monitoring applications, edge AI enables wearable devices to analyze user physiological data in real-time, providing real-time health feedback to users without uploading data to the ”cloud” for processing [77]. These devices harness DL technologies to scrutinize user health metrics and furnish guidance for subsequent lifestyle modifications, yielding substantial benefits, particularly for infants, young children, and the elderly. Due to data not being uploaded to the cloud, it can effectively protect user privacy and security. In addition, in the field of intelligent transportation, AI technology provides timely guidance for vehicle operation by analyzing local data. Similarly, many frameworks utilize DL techniques to predict parking occupancy rates, reduce the time required for vehicles to search for parking spaces, and improve urban traffic management [6]. These user cases demonstrate the enormous potential of edge AI technology in improving the performance of mobile IoT applications and enhancing user experience.
@@ V-D1 Heuristic
People often refer to methods inspired by the laws of nature or the experiences and rules of specific problems as heuristic algorithms. The current heuristic algorithms are not entirely based on natural laws, but also come from human accumulated work experience. They supply a practical solution for each instance of the combinatorial problem to be solved and maintaining an acceptable cost in terms of computation time and space, and the degree of deviation between the practical solution and the optimal solution may not be ascertainable before the implementation in advance [60]. Heuristic algorithms are a technique that enables the search for the best possible solution within an acceptable computational cost, but may not necessarily guarantee that the obtained solution is the best scheme. In most cases, it is impossible to describe the degree of approximation between the obtained solution and the optimal solution.
@@ V-D3 Machine Learning
As is well known, ML technique forms the foundation of AI and plays a pivotal role in many applications, such as recommendation systems, text generation, and so on. As edge AI gains prominence, ML is increasingly finding its way onto devices located on the fringes of the network. These edge devices prioritize data processing close to its source, aiming to minimize transmission delays, accelerate response times, and alleviate the reliance on centralized servers [6]. Consequently, edge AI systems are often tasked with handling vast volumes of data in real-time scenarios. Following certain ML algorithms, such as sophisticated DL models, these systems can swiftly analyze and process data at the edge, extract crucial insights, and empower a diverse array of real-time applications encompassing autonomous vehicle operations, intelligent manufacturing processes, and robust security systems. In addition, ML techniques can be employed to detect anomalies in real-time on edge devices, such as detecting product quality issues during the manufacturing process, or detecting changes in patient health status in the medical field [81].
@@ V-E Resource Management
Edge computing, which deploys computing, storage, network, and other resources at the network’s edge, can drastically minimize data transmission delay, enhance data processing efficiency, and relieve bandwidth demand on the core network [4]. However, with the growing number of edge devices and the complexity of applications, how to efficiently manage these resources has become an urgent problem that must be addressed. This section will review how to make resource management in edge environment from the aspects of resource provisioning, resource allocation, application placement, and workload distribution and prediction.
@@ V-E4 Workload Distribution and Prediction
Furthermore, accurate predictions can lead to more rational resource management. This is beneficial for both users and service providers. Previous research and technologies such as ARIMA [88] and Holt Winter[89] were based on linear temporal prediction. However, these technologies often exhibit poor prediction accuracy. With the rise of DL, technologies such as neural networks have been widely applied in data prediction, such as weather, transportation, and finance systems. Especially for recurrent neural networks (RNN), their inputs not only focus on current data, but also contain information from a period of time in the past. Therefore, it is often used to predict time-series related data. Using recurrent neural networks to predict workload in edge computing is a promising approach worthy of consideration.
@@ V-F1 Reduced
The training cost and efficiency of AI models are important metrics to assess the quality of the model. Nowadays, many AI giants are progressively increasing the size of model parameters and the volume of training data. The model parameters of GPT-3.5 have reached 175 billion [91]. While this approach significantly enhances model accuracy, it also markedly escalates training costs and hardware requirements, necessitating a trade-off between accuracy and cost. In this way, we need to strike a balance between accuracy and cost. In recent years, there have been numerous studies in this area, such as DenseNet [92], EfficientNet [93] and EfficientNetV2 [94]. The goal of these works is to train models to achieve satisfactory accuracy with fewer model parameters. In the field of edge computing, where hardware resources are limited, the reduced size model will certainly become an important trend of edge AI in the future.
@@ V-G1 Computational
For many microservices, such as the web services mentioned earlier, their performance bottleneck often is not in CPU but in disk read and write speed, as most of the time is spent accessing databases. In other words, they are IO intensive services rather than computationally intensive. In order to reduce network latency, microservices are usually deployed in edge nodes. How to reduce the performance loss caused by slow disk read and write speeds is a problem we need to consider.
V-G2 Hardware
@@
V-G3 Platform
Due to the rise of edge computing, the world’s major technology giants have also launched their own edge computing platforms. For example, Amazon’s AWS IoT Greengrass, Microsoft’s Azure IoT Edge, and Google’s Cloud IoT Edge. They all support the effective operation of AI models on edge devices, providing service management and data analysis capabilities. Other open source platforms also deserve attention, such as KubeEdge and OpenYurt, which are extensions of Kubernetes in the field of edge computing and provide container management, automatic operation and maintenance and other functions.
@@ V-H3 Network
Distributed Denial of Service (DDoS) attack, which causes significant economic losses to society every year, is one of the most common attack methods in computer networks, and it also has strong destructive power on IoT devices. From the time of the 2016 botnet Mirai attack on KrebsOnSecurity [106] and Dyn [107], it can be seen that DDoS attacks are seriously threatening the security of IoT applications. With the development of edge computing, the threat of such attacks to large-scale IoT devices is growing, which may lead to incalculable economic losses. For example, in the field of automation, AI technology is widely used to make decisions and adjust plans. If edge AI is subjected to network attacks, it can lead to AI models making incorrect decisions, resulting in product quality issues.
Although edge nodes exhibit the potential to isolate most of the IoT data at the network edge and detect and intercept attacks near the source in the first place, they encounter significant challenges in practical applications. The main reason is that edge nodes are unable to capture the aggregated network traffic required for IoT DDoS detection, nor can they scale and provide the necessary resources like elastic clouds [108]. Therefore, directly deploying existing cloud based defense solutions on edge nodes is far from achieving ideal results. We need to redesign the DDoS defense scheme based on edge computing to solve the special and severe security problems in the edge environment.
@@ V-I1 Container
Containers are a software virtualization technology that laid the foundation for the development of microservices. Nowadays, containers are also widely used in the field of edge computing.There are also many researches on containers in edge computing, which stems from the growing demand of users for millisecond delay computing. In  [109], the authors elucidate the concepts of container placement and migration between edge servers, and proposes a container scheduling framework grounded in multi-objective optimization models or graph network models.
In addition, some open-source container orchestration and scheduling frameworks are worth paying attention to, such as KubeEdge. KubeEdge can extend its powerful cloud computing capabilities to edge devices. Especially suitable for some AI applications, model training can be completed in the cloud and then deployed to edge devices. In addition, KubeEdge can optimize scheduling performance based on different AI application business scenarios by configuring the algorithm and parameters of kube-scheduler.
@@ V-I2 Task
In the scheduling task of edge computing, there are usually two problems to be solved: scheduling time and resource allocation. Most previous research [110, 111, 112] on task scheduling has focused on these two aspects. With the advancement of AI technology, ML technology has shown unique advantages in task scheduling. Markov Decision Process (MDP) is an robust and effective method for modeling temporal data and providing high-precision predictions. The problem of resource allocation in edge devices can be described as MDP, and the deep Q network (DQN) algorithm uses multiple replay memories to minimize the total delay and resource utilization. the study in [113] addresses the intricate issue of joint task offloading and resource allocation problem for computationally intensive tasks in fog computing. This intricate problem is formulated as a partially observable MDP, and the Deep Recursive Q-Network (DRQN) algorithm is adopted to approximate the optimal value function.
@@ V-I3 Pod
In container orchestration systems such as Kubernetes, Pod is the smallest unit of work composed of several containers. Pod scheduling is the process of assigning Pods to a node based on a certain algorithm strategy, which is of great significance for ensuring high availability, resource utilization, and performance of the system. Pod scheduling is mainly controlled by kube-scheduler, and its process includes two stages: screening and scoring [114]. During the filtering phase, the scheduler checks all nodes to determine which ones have the resources (such as CPU and memory) and other requirements (such as node selector labels) needed to run Pod. Then, the selected node will enter the scoring stage, and the scheduler will rate each node based on a series of criteria such as node affinity, resource utilization, etc. The node scoring highest will be designated as the running location for Pod. kube-scheduler supports custom scheduling plugins, and users can develop some extension plugins based on the business characteristics of the enterprise.
@@ V-J1 Stateful vs Stateless containers
Stateful container: The so-called state essentially refers to the data in the running container. When migrating such containers, it is usually necessary to migrate their data together, such as a database. Due to the involvement of data replication, such containers need to consider issues such as data loss and data integrity. Specific migration tools or strategies may be needed to ensure accurate migration and recovery of data [115]. All containers managed by a StatefulSet controller in Kubernetes are considered as stateful.
Stateless containers: These containers are containers that do not save any state during runtime. For example, a web server that provides services for static pages, treats each request as independent, and the container does not need to remember previous interactions. The migration process is very simple, just pull up the container on other nodes and delete the container from the original node. All Pods under a Deployment in Kubernetes are stateless.
@@ V-J2 Inter versus Intra cluster migrations
Inter-cluster migration: When a company or organization needs to migrate its data center from one geographic location to another, inter-cluster migration is an indispensable step. Inter-cluster migrations involve node migration between different clusters, typically requiring consideration of cross-cluster communication factors such as network latency and bandwidth limitations [116]. Due to the collaborative work of multiple clusters and nodes involved in cross-cluster migration, the migration process is relatively complex and requires ensuring data consistency and service continuity.
Intra-cluster migration: In a cluster, migration within the cluster can take effect when a node experiences performance degradation or longer response time due to excessive workload. Administrators or automation tools can migrate a portion of the workload (such as containers, virtual machines, or services) on that node to other nodes in the cluster to balance the load and optimize performance [42]. Compared to inter-cluster migrations, The complexity of intra-cluster migrations is relatively low because it only involves nodes and data migration within the same cluster.
@@ V-J3 Migrations at cloud/edge/fog
Edge computing has become an important technical support for the development of IoT [118]. A thorny problem in edge computing is service migration, especially in the mobile IoT device environment. Due to the limited coverage of a single edge server network, the migration of mobile services between servers is likely to reduce the QoS of the services. State preservation of services (such as stateful services), data loss, and cost control have become challenges in the migration of edge computing services.
Migration at fog is the process of migrating applications, services, or data from traditional centralized data centers or cloud environments to a fog computing environment. The purpose of this migration is to achieve low latency, bandwidth optimization, enhanced security, improved scalability, and fault tolerance. Fog migration involves redesigning applications to adapt to the distributed architecture of fog computing, including modular design and the ability to handle network dynamics [45, 119]. Fog migration can provide more effective support for IoT devices, mobile devices and other applications that need rapid response, and achieve the goal of intelligent edge computing.
@@ V-J4 Simulations versus real-world testbed migrations
Simulations: It uses models to replace actual or conceptual systems for training, analysis, argumentation, experimentation, experimentation, and planning methods, techniques, and activities [4]. Simulations can predict system performance and efficiency, validate and iterate modeling and simulation through real experimental data, support, optimize and expand experimental identification, accelerate development and reduce risk costs [60].
Real world testbed migrations: Its definition is the process of testing and validating a system or application in a real physical environment, involving the migration of the system or application from one environment to another. Since it is conducted in a real-world environment, it can directly evaluate the performance, reliability, and safety of the system or application under actual operating conditions to ensure that it meets practical needs [1].
When conducting container migration testing and validation, simulations and real-world testing platforms are usually combined. Simulation can quickly validate concepts and strategies in the early stages, while real-world testing platforms are used to test and optimize migration strategies under conditions close to actual operational environments. This combination of methods can balance cost, time, and accuracy of results, providing a comprehensive evaluation for fog migration.

### v2

@@ V-A3 Edge
The concept of edge computing is relative to cloud computing. The processing method of cloud computing is to upload all data to the cloud data center or server in the computing resource set for processing. Any request to access this information must be submitted to the cloud for processing [35]. Edge computing is a computing model that moves resources and provides edge intelligent services at the network edge near objects or data sources, which improves service quality and information security. In brief, edge computing analyzes the data produced by the terminal directly in the local device or network near the data generation, without transmitting the data to the cloud data processing center [68]. Compared with cloud computing, edge computing has shorter network latency, less resource consumption, and higher security of the data to the cloud data processing center.
The so-called edge AI combines two emerging technologies: edge computing and AI. However, the implementation of edge computing is based on the same basic premise, that is, generating, collecting, storing, processing and managing data locally rather than in remote data centers. Edge AI improves this concept to the device level, using ML to imitate human reasoning to reach user interaction points, such as computers, edge servers or IoT devices [69]. Typically, these devices can operate without an Internet connection and make decisions independently. Well-known examples of edge AI technologies include virtual assistants, such as GPT-4o, Apple’s Siri, or Amazon Alexa. When the user says “hey”, these tools will recognize and learn what the user is saying (i.e. ML), interact with cloud-based application programming interfaces (APIs), and store the learned knowledge locally.
@@ V-C2 Mobile
The emergence of emerging AI such as DL has brought innovations to mobile animal networking. In the field of mobile IoT, edge AI achieves real-time and efficient data processing by deploying the computing power of AI at the edge of devices. For example, in smart health monitoring applications, edge AI enables wearable devices to analyze user physiological data in real-time, providing real-time health feedback to users without uploading data to the ”cloud” for processing [81]. These devices harness DL technologies to scrutinize user health metrics and furnish guidance for subsequent lifestyle modifications, yielding substantial benefits, particularly for infants, young children, and the elderly. Due to data not being uploaded to the cloud, it can effectively protect user privacy and security. In addition, in the field of intelligent transportation, AI technology provides timely guidance for vehicle operation by analyzing local data. Similarly, many frameworks utilize DL techniques to predict parking occupancy rates, reduce the time required for vehicles to search for parking spaces, and improve urban traffic management [6]. These user cases demonstrate the enormous potential of edge AI technology in improving the performance of mobile IoT applications and enhancing user experience.
@@ V-D1 Heuristic
People often refer to methods inspired by the laws of nature or the experiences and rules of specific problems as heuristic algorithms. The current heuristic algorithms are not entirely based on natural laws, but also come from human accumulated work experience. They supply a practical solution for each instance of the combinatorial problem to be solved and maintain an acceptable cost in terms of computation time and space, and the degree of deviation between the practical solution and the optimal solution may not be ascertainable before the implementation in advance [31]. Heuristic algorithms are a technique that enables the search for the best possible solution within an acceptable computational cost, but may not necessarily guarantee that the obtained solution is the best scheme. In most cases, it is impossible to describe the degree of approximation between the obtained solution and the optimal solution.
@@ V-D3 Machine Learning
As is well known, the ML technique forms the foundation of AI and plays a pivotal role in many applications, such as recommendation systems, text generation, and so on. As edge AI gains prominence, ML is increasingly finding its way onto devices located on the fringes of the network. These edge devices prioritize data processing close to its source, aiming to minimize transmission delays, accelerate response times, and alleviate the reliance on centralized servers [6]. Consequently, edge AI systems are often tasked with handling vast volumes of data in real-time scenarios. Following certain ML algorithms, such as sophisticated DL models, these systems can swiftly analyze and process data at the edge, extract crucial insights, and empower a diverse array of real-time applications encompassing autonomous vehicle operations, intelligent manufacturing processes, and robust security systems. In addition, ML techniques can be employed to detect anomalies in real-time on edge devices, such as detecting product quality issues during the manufacturing process, or detecting changes in patient health status in the medical field [85].
@@ V-E Resource Management
Edge computing, which deploys computing, storage, network, and other resources at the network’s edge, can drastically minimize data transmission delay, enhance data processing efficiency, and relieve bandwidth demand on the core network [4]. However, with the growing number of edge devices and the complexity of applications, how to efficiently manage these resources has become an urgent problem that must be addressed. This section will review how to make resource management in an edge environment from the aspects of resource provisioning, resource allocation, application placement, and workload distribution and prediction.
@@ V-E4 Workload Distribution and Prediction
Furthermore, accurate predictions can lead to more rational resource management. This is beneficial for both users and service providers. Previous research and technologies such as ARIMA [92] and Holt Winter[93] were based on linear temporal prediction. However, these technologies often exhibit poor prediction accuracy. With the rise of DL, technologies such as neural networks have been widely applied in data prediction, such as weather, transportation, and finance systems. Especially for recurrent neural networks (RNNs), their inputs not only focus on current data, but also contain information from a period of time in the past. Therefore, it is often used to predict time-series related data. Using RNNs to predict workload in edge computing is a promising approach worthy of consideration.
@@ V-F1 Reduced
The training cost and efficiency of AI models are important metrics to assess the quality of the model. Nowadays, many AI giants are progressively increasing the size of model parameters and the volume of training data. The model parameters of GPT-3.5 have reached 175 billion [95]. While this approach significantly enhances model accuracy, it also markedly escalates training costs and hardware requirements, necessitating a trade-off between accuracy and cost. In this way, we need to strike a balance between accuracy and cost. In recent years, there have been numerous studies in this area, such as DenseNet [96], EfficientNet [97] and EfficientNetV2 [98]. The goal of these works is to train models to achieve satisfactory accuracy with fewer model parameters. In the field of edge computing, where hardware resources are limited, the reduced-size model will certainly become an important trend of edge AI in the future.
@@ V-G1 Computational Heterogeneity
For many microservices, such as the web services mentioned earlier, their performance bottleneck often is not in CPU but in disk read and write speed, as most of the time is spent accessing databases. In other words, they are IO-intensive services rather than computationally intensive. In order to reduce network latency, microservices are usually deployed in edge nodes. How to reduce the performance loss caused by slow disk read and write speeds is a problem we need to consider.
V-G2 Hardware Heterogeneity
@@
V-G3 Platform Heterogeneity
Due to the rise of edge computing, the world’s major technology giants have also launched their own edge computing platforms. For example, Amazon’s AWS IoT Greengrass, Microsoft’s Azure IoT Edge, and Google’s Cloud IoT Edge. They all support the effective operation of AI models on edge devices, providing service management and data analysis capabilities. Other open-source platforms also deserve attention, such as KubeEdge and OpenYurt, which are extensions of Kubernetes in the field of edge computing and provide container management, automatic operation and maintenance and other functions.
@@ V-H3 Network
Distributed Denial of Service (DDoS) attack, which causes significant economic losses to society every year, is one of the most common attack methods in computer networks, and it also has strong destructive power on IoT devices. From the time of the 2016 botnet Mirai attack on KrebsOnSecurity [110] and Dyn [111], it can be seen that DDoS attacks are seriously threatening the security of IoT applications. With the development of edge computing, the threat of such attacks on large-scale IoT devices is growing, which may lead to incalculable economic losses. For example, in the field of automation, AI technology is widely used to make decisions and adjust plans. If edge AI is subjected to network attacks, it can lead to AI models making incorrect decisions, resulting in product quality issues.
Although edge nodes exhibit the potential to isolate most of the IoT data at the network edge and detect and intercept attacks near the source in the first place, they encounter significant challenges in practical applications. The main reason is that edge nodes are unable to capture the aggregated network traffic required for IoT DDoS detection, nor can they scale and provide the necessary resources like elastic clouds [112]. Therefore, directly deploying existing cloud-based defense solutions on edge nodes is far from achieving ideal results. We need to redesign the DDoS defense scheme based on edge computing to solve the special and severe security problems in the edge environment.
@@ V-I1 Container
Containers are a software virtualization technology that laid the foundation for the development of microservices. Nowadays, containers are also widely used in the field of edge computing. There is also much research on containers in edge computing, which stems from the growing demand of users for millisecond delay computing. In  [113], the authors elucidate the concepts of container placement and migration between edge servers, and propose a container scheduling framework grounded in multi-objective optimization models or graph network models.
In addition, some open-source container orchestration and scheduling frameworks are worth paying attention to, such as KubeEdge. KubeEdge can extend its powerful cloud computing capabilities to edge devices. Especially suitable for some AI applications, model training can be completed in the cloud and then deployed to edge devices. In addition, KubeEdge can optimize scheduling performance based on different AI application business scenarios by configuring the algorithm and parameters of the kube-scheduler.
@@ V-I2 Task
In the scheduling task of edge computing, there are usually two problems to be solved: scheduling time and resource allocation. Most previous research [114, 115, 116] on task scheduling has focused on these two aspects. With the advancement of AI technology, ML technology has shown unique advantages in task scheduling. Markov Decision Process (MDP) is a robust and effective method for modeling temporal data and providing high-precision predictions. The problem of resource allocation in edge devices can be described as MDP, and the deep Q network (DQN) algorithm uses multiple replay memories to minimize the total delay and resource utilization. the study in [117] addresses the intricate issue of joint task offloading and resource allocation problems for computationally intensive tasks in fog computing. This intricate problem is formulated as a partially observable MDP, and the Deep Recursive Q-Network (DRQN) algorithm is adopted to approximate the optimal value function.
@@ V-I3 Pod
In container orchestration systems such as Kubernetes, Pod is the smallest unit of work composed of several containers. Pod scheduling is the process of assigning Pods to a node based on a certain algorithm strategy, which is of great significance for ensuring high availability, resource utilization, and performance of the system. Pod scheduling is mainly controlled by kube-scheduler, and its process includes two stages: screening and scoring [118]. During the filtering phase, the scheduler checks all nodes to determine which ones have the resources (such as CPU and memory) and other requirements (such as node selector labels) needed to run Pod. Then, the selected node will enter the scoring stage, and the scheduler will rate each node based on a series of criteria such as node affinity, resource utilization, etc. The node scoring the highest will be designated as the running location for Pod. kube-scheduler supports custom scheduling plugins, and users can develop some extension plugins based on the business characteristics of the enterprise.
@@ V-J1 Stateful vs Stateless containers
(i) Stateful containers: The so-called state essentially refers to the data in the running container. When migrating such containers, it is usually necessary to migrate their data together, such as a database. Due to the involvement of data replication, such containers need to consider issues such as data loss and data integrity. Specific migration tools or strategies may be needed to ensure accurate migration and recovery of data [119]. All containers managed by a StatefulSet controller in Kubernetes are considered stateful.
(ii) Stateless containers: These containers are containers that do not save any state during runtime. For example, a web server that provides services for static pages, treats each request as independent, and the container does not need to remember previous interactions. The migration process is very simple, just pull up the container on other nodes and delete the container from the original node. All Pods under a Deployment in Kubernetes are stateless.
@@ V-J2 Inter versus Intra cluster migrations
Inter-cluster and intra-cluster migration are discussed briefly below:
(i) Inter-cluster migration: When a company or organization needs to migrate its data center from one geographic location to another, inter-cluster migration is an indispensable step. Inter-cluster migrations involve node migration between different clusters, typically requiring consideration of cross-cluster communication factors such as network latency and bandwidth limitations [120]. Due to the collaborative work of multiple clusters and nodes involved in cross-cluster migration, the migration process is relatively complex and requires ensuring data consistency and service continuity.
(ii) Intra-cluster migration: In a cluster, migration within the cluster can take effect when a node experiences performance degradation or longer response time due to excessive workload. Administrators or automation tools can migrate a portion of the workload (such as containers, virtual machines, or services) on that node to other nodes in the cluster to balance the load and optimize performance [63]. Compared to inter-cluster migrations, The complexity of intra-cluster migrations is relatively low because it only involves nodes and data migration within the same cluster.
@@ V-J3 Migrations at cloud/edge/fog
Edge computing has become an important technical support for the development of IoT [122]. A thorny problem in edge computing is service migration, especially in the mobile IoT device environment. Due to the limited coverage of a single-edge server network, the migration of mobile services between servers is likely to reduce the QoS of the services. State preservation of services (such as stateful services), data loss, and cost control have become challenges in the migration of edge computing services.
Migration at fog is the process of migrating applications, services, or data from traditional centralized data centers or cloud environments to fog computing environments. The purpose of this migration is to achieve low latency, bandwidth optimization, enhanced security, improved scalability, and fault tolerance. Fog migration involves redesigning applications to adapt to the distributed architecture of fog computing, including modular design and the ability to handle network dynamics [66, 123]. Fog migration can provide more effective support for IoT devices, mobile devices and other applications that need rapid response, and achieve the goal of intelligent edge computing.
@@ V-J4 Simulations versus real-world testbed migrations
(i) Simulations: It uses models to replace actual or conceptual systems for training, analysis, argumentation, experimentation, experimentation, and planning methods, techniques, and activities [4]. Simulations can predict system performance and efficiency, validate and iterate modeling and simulation through real experimental data, support, optimize and expand experimental identification, accelerate development and reduce risk costs [31].
(ii) Real-world testbed migrations: Its definition is the process of testing and validating a system or application in a real physical environment, involving the migration of the system or application from one environment to another. Since it is conducted in a real-world environment, it can directly evaluate the performance, reliability, and safety of the system or application under actual operating conditions to ensure that it meets practical needs [1].
When conducting container migration testing and validation, simulations and real-world testing platforms are usually combined. Simulation can quickly validate concepts and strategies in the early stages, while real-world testing platforms are used to test and optimize migration strategies under conditions close to actual operational environments. This combination of methods can balance the cost, time, and accuracy of results, providing a comprehensive evaluation for fog migration.